# Mind the Gap. Doubling Constant Parametrization of Weighted Problems: TSP, Max-Cut, and More

### v1

@@ 1 Introduction
Indeed, there are many problems that resist improvements over the standard algorithm. Yet, we have arguments to be optimistic: The amount of work in this direction has led to impressive speedups for well-studied problems such as Max-Cut [maxcut-williams], Hamiltonicity [hamiltonicity, below_all_subsets1], Sched [below_all_subsets2], Independent Set [below_all_subsets3], Dominating Set [below_all_subsets3, below_all_subsets4], and Min $k$-Cut [min_k_cut]. To some extent, these speedups have been achieved mainly for what we will refer to as unweighted problems. The weighted versions of many of these problems are still stuck with the textbook solutions that accept only modest, mostly polynomial-time, improvements.
A concrete example is that of the Traveling Salesman Problem (TSP), which asks to find the shortest tour through $n$ cities. The textbook algorithm in $O(2^{n}n^{2})$-time remains, up to some modest polynomial improvements [chan_shaving_logs], the best up to date for the general setting [bellman_dp, held_karp_dp, nederlof_tsp]. In contrast, its unweighted counterpart, the Hamiltonicity problem, has been (non-trivially) sped up to $O^{*}(1.66^{n})$-time by Björklund [hamiltonicity]. [^1] The $O^{*}$-notation hides polynomial factors in the input size.
Interestingly, virtually every new result on the unweighted version of a weighted problem has its own theorem stating that the original problem can be solved in the time of its unweighted counterpart, yet with an extra multiplicative pseudo-polynomial dependence on the largest input weight; this is known as the “(polynomial) embedding technique”. Notably, this situation also concerns TSP: Björklund proved that Hamiltonicity can be leveraged to solve TSP by paying the price of a near-linear time-factor depending on the largest edge weight $W$ [hamiltonicity, Thm. 3]. This becomes impractical for arbitrarily large $W$. The same holds true for the weighted Max-Cut problem: its unweighted counterpart can be solved in $O^{*}(2^{\omega n/3})$-time [maxcut-williams, koivisto-correct], where $\omega<2.371339$ is the matrix multiplication exponent [curr-omega], while its weighted version has to “accept” the additional pseudo-polynomial factor. On the other hand, if the weights are polynomially bounded, then the problematic factor becomes polynomial, and the running time of the weighted version matches, in the $O^{*}$-sense, that of the unweighted version.
@@
Given the ubiquity of the embedding technique, it is natural to ask whether the gap between the running times of the weighted and unweighted versions of a problem is overcome only in the regime of polynomially bounded weights. A negative answer would be consistent with the widespread adoption of the embedding approach, with its inherent pseudo-polynomial dependence [maxcut-williams, koivisto-correct, fsc, set_partitioning, hamiltonicity, saving_space, Cygan_book]. In contrast, a positive answer would re-open the door to a more systematic re-examination of the weighted vs. unweighted gap, in the spirit of the “Losing Weights by Gaining Edges” technique for node-weighted problems [losing-ws] or in terms of approximation guarantees [apx-ws].
@@ 1.1 Our Results
To maintain notation consistency with prior work on unweighted problems [esa-paper], we prefix problem names with $\mathcal{C}$-” if the set of input weights has doubling constant $\mathcal{C}$. We also write $O_{\mathcal{C}}$ to note that we suppressed factors that depend only on $\mathcal{C}$. Under this notation, all of our results conform to the following template: {restatable}[]corollaryDCTSP If Hamiltonicity can be solved in time $T(n)$, then $\mathcal{C}$-TSP can be solved in time $O_{\mathcal{C}}^{*}(T(n))$. The applicability of our work goes beyond TSP. Indeed, we can consider any weighted NP-hard problem that operates over the $(\min,+)$ and $(\max,+)$ semirings, such as the Weighted Max-Cut, Edge-Weighted $k$-Clique, or Minimum Steiner Tree problems. [^2] We kindly refer the reader to the relevant section for the problem definitions. The only requirement is that the problem at hand satisfies a property $\phi$: its objective value is an additive combination of input weights, and it admits an (algebraic) algorithm as seen earlier; see Prop. 1 for a formal treatment.
@@
[Meta-algorithm]theoremMETATHM If problem $\mathbf{P}_{w}$ satisfies property $\phi$ and the unweighted version can be solved in time $O(T(n))$, then $\mathcal{C}$-$\mathbf{P}_{w}$ can be solved in time $O_{\mathcal{C}}^{*}(T(n))$.
@@ 1.2 Overview of Approach
Let $A$ denote the set of input weights of the problem at hand, such as the edge weights in TSP. The first observation is that the objective value of the problem is contained in a $q$-fold sumset of $A$, i.e., $A+\ldots+A$ with $A$ occurring $q$ times, where $q$ is problem-specific; for TSP we have $q=n$. The second observation is that if $A$ has small doubling, then the cardinality of this very $q$-fold sumset is polynomial w.r.t. $O_{\mathcal{C}}$. [^3] Noteworthy, Plünnecke’s inequality is too loose in this setting, leading to an exponential bound. This suggests that, in principle, we could run the bounded-input algorithm. However, the values may still be numerically as large as $\Theta(q\cdot\max A)$. The missing piece is a decomposition of the weights into constant-size coefficient-tuples coming from the generalized arithmetic progression $A$ is a subset of (cf. Freiman’s theorem [freiman1]).
@@ 1.3 Related Work
Our work continues the effort to consider important results from additive combinatorics [tao_vu] as a new toolbox for algorithm design, a connection highlighted in works such as Ref. [trevisan_ac, ac_cs_view3, ac_cs_view1, ac_cs_view2]. One of the first seminal results is that of Chan and Lewenstein [chan-lewenstein], who used the (constructive) Balog-Szemerédi-Gowers (BSG) theorem to speed up the monotone variant of the min-plus convolution on small integers. Recently, this has been further used and extended in #APSP [fredman_meets_domi], 3SUM [3sum_sidon], listing 4-cycles [abf], and approximate distance oracles [abf]. We observe that the constructive version of a fundamental theorem in additive combinatorics bears fruit in algorithm design. To this end, Randolph and Węgrzycki [esa-paper] made Freiman’s theorem [freiman1], another fundamental result in additive combinatorics, constructive. Their main applications were to SubsetSum and ILP-Feasibility, which have an inherent additive structure. Indeed, we extend their results to the area of weighted problems.
@@
Repurposing the unweighted algorithm for the weighted version of a problem is a common theme in work that improves over the textbook solution, as in seminal results such as Ref. [maxcut-williams, fsc, set_partitioning, hamiltonicity]; notably, in all these examples, the weighted problem operates over the $(\min,+)$ and $(\max,+)$ semirings. The key idea is to represent the input weights of the problem instance as monomials so that the bounded-input algorithm can work in the $(+,\times)$ ring. The drawback is that this introduces a pseudo-polynomial dependence in the running time, since operations now incur an additional $\widetilde{O}(W)$ factor, [^4] The $\widetilde{O}-$notation hides polylogarithmic factors in the input size. where $W$ is the largest input weight. For instance, undirected Hamiltonicity runs in $O^{*}(1.66^{n})$-time, hence TSP can be solved in $O^{*}(1.66^{n}W)$-time. This, however, becomes impractical for arbitrarily large $W$. Still, if $W$ is polynomially bounded, then up to polynomial factors, the weighted problem becomes equivalent to the unweighted one. For example, if $W=\text{poly}(n)$, then TSP can be solved in time $O^{*}(1.66^{n})$, matching its unweighted counterpart.
@@
The pseudo-polynomial factor is often unwanted. One way to shave it is to enter the realm of approximation. In fact, there is a simple exercise to turn any bounded-input algorithm for a problem into an exponential-time $(1+\varepsilon)$-approximation algorithm. For the Max-Cut problem, Williams mentions this idea in passing [maxcut-williams], inspired by the literature on the all-pairs shortest paths problem [zwick_apsp]; see, e.g., Ref. [approx-min-sum-conv] for more such examples. It is interesting to ask whether the doubling constant could play a role in approximation algorithms.
@@ 2 Preliminaries
As we will only deal with exponential-time algorithms, we use the $O^{*}$-notation to hide polynomial factors in the input size. For example, $O^{*}(2^{n})=O(2^{n}\text{poly}(n))$. In particular, this notation will never hide pseudo-polynomial factors depending on the input weights; specifically, we mean the largest input weight, which will be denoted by $W$. Furthermore, we use the $O_{\mathcal{C}}$-notation, already established in Ref. [esa-paper], to denote that we suppressed factors depending on $\mathcal{C}$, the doubling constant on which we parameterize the time-complexities. For instance, $O_{\mathcal{C}}(2^{n})=f(\mathcal{C})\cdot O(2^{n})$ for a computable function $f$. To put into context, $O_{\mathcal{C}}^{*}(2^{n})$ will thus mean $f(\mathcal{C})\cdot O^{*}(2^{n})$.
@@ 2.1 Embedding Technique
The solution polynomial abstracts away whether the problem is a minimization or maximization problem: for minimization we select the monomial with the smallest exponent, whereas for maximization we select the monomial with the largest exponent. To this end, we refer to an algorithm as algebraic, if it returns a solution polynomial for an instance of a problem. [^6] An alternative attribute could be non-combinatorial. For a treatment on the definition of a combinatorial algorithm, see Ref. [fast-bmm].
@@ 2.2 Additive Combinatorics
Given a set $A$ of small doubling, Freiman’s theorem [freiman1] tells that, indeed, $A$ is included in a GAP $G$, whose dimension is independent on $|A|$. In this work, we will need the constructive Freiman’s theorem, recently designed by Randolph and Węgrzycki [esa-paper]:
Theorem 2.2 (Constructive Freiman’s Theorem [esa-paper]).
@@
The above theorem gives us a GAP of dimension $d(\mathcal{C})=2^{\mathcal{C}^{O(1)}}$ and volume $v(\mathcal{C})=2^{2^{\mathcal{C}^{O(1)}}}n$, as in the original Freiman’s theorem [freiman1]. Noteworthy, later works do optimize $d$ and $v$ [freiman-better-1, freiman-better-2, freiman-better-3, freiman-better-4], yet these improvements are not captured by the constructive version.
@@ 4 Applications
In the following, we review several weighted NP-hard problems and show how to solve their small-doubling instances faster than the traditional algorithm by leveraging the previous meta-algorithm. This enables running times in the regime of the unweighted version of the problem for a large suite of problems, such as TSP, Weighted Max-Cut, Edge-Weighted $k$-Clique, and Minimum Steiner Tree. We stress that, compared to prior work [esa-paper], we consider weighted problems that operate over tropical semirings.
@@ 4.1 TSP Meets Small Doubling
Condition 1 holds since a tour’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as Björklund’s algorithm [hamiltonicity] for undirected Hamiltonicity, which runs in time $O^{*}(1.66^{n})$, and solves the weighted case in time $O^{*}(1.66^{n}\,W)$ [hamiltonicity, Thm. 3]. Due to the algorithm’s design, namely that it counts cycle covers of $n$ arcs, we have $\lambda\leq n$. ∎
@@ 4.2 Weighted Max-Cut Under Small Doubling
Another well-studied problem whose unweighted version has been sped up and fits the small doubling setting well is the Weighted Max-Cut problem. In a breakthrough result, Williams [maxcut-williams] improved the time complexity of the unweighted Max-Cut to $O^{*}(2^{\omega n/3})$. We show that we can leverage this result to the weighted setting when the edge weights form an integer set that has small doubling. We formally define the doubling constant parametrization of the problem in the following:
@@
Condition 1 holds since a cut’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as Williams’s algebraic algorithm [maxcut-williams] (and later improved by Koivisto [koivisto-correct]) for the unweighted Max-Cut problem, which runs in time $O^{*}(2^{\omega n/3})$, and solves the weighted case in time $O^{*}(2^{\omega n/3}\,W)$ [koivisto-correct, Thm. 3]. ∎
@@ 4.3 Small Doubling and Edge-Weighted $k$-Cliques
We now turn to a problem whose complexity is a rather popular hypothesis in hardness proofs [clique-hyp-1, clique-hyp-2, clique-hyp-3]: the edge-weighted $k$-clique problem. Compared to the problems considered so far, the algebraic algorithm for the unweighted $k$-Clique has already been known from the 80’s [uw-k-clique]. We show that it can be leveraged in the doubling constant parametrization of the edge-weighted version. Note that, apart from a mention in passing in Lincoln et al. [clique-hyp-3] that Nešetřil-Poljak’s algorithm can also be used for polynomially bounded weights, we could not find any reference to a proof. For completeness, we show this in Lemma B (Appendix B).
@@
Condition 1 holds since a clique’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as Nešetřil-Poljak’s algebraic algorithm [uw-k-clique] for the unweighted version of the problem (or the improvement by Eisenbrand and Grandoni [faster-uw-k-clique] for some values of $k$ using rectangular matrix multiplication), which runs in time $O^{*}(n^{k\omega/3})$, and solves the weighted case in time $O^{*}(n^{k\omega/3}\,W)$ (Lemma B). Notably, due to its design, namely that each original input weight is doubled in the auxiliary graph, the bounded-input algorithm (Lemma B) has a $\lambda\leq 2|w(I)|$. ∎

### v2

# Mihail Stoian

@@ 1 Introduction
Indeed, there are many problems that resist improvements over the standard algorithm. Yet, we have arguments to be optimistic: The amount of work in this direction has led to impressive speedups for well-studied problems such as Max-Cut [Wil05], Hamiltonicity [Bjö10, Bjö16], Sched [CPPW14], Independent Set [FGK09], Dominating Set [FGK09, vRNvD09], and Min $k$-Cut [LSS23]. To some extent, these speedups have been achieved mainly for what we will refer to as unweighted problems. The weighted versions of many of these problems are still stuck with the textbook solutions that accept only modest, mostly polynomial-time, improvements.
A concrete example is that of the Traveling Salesman Problem (TSP), which asks to find the shortest tour through $n$ cities. The textbook algorithm in $O(2^{n}n^{2})$-time remains, up to some modest polynomial improvements [Cha13], the best up to date for the general setting [Bel62, HK61, Ned20]. In contrast, its unweighted counterpart, the Hamiltonicity problem, has been (non-trivially) sped up to $O^{*}(1.66^{n})$-time by Björklund [Bjö10]. [^2] The $O^{*}$-notation hides polynomial factors in the input size.
Interestingly, virtually every new result on the unweighted version of a weighted problem has its own theorem stating that the original problem can be solved in the time of its unweighted counterpart, yet with an extra multiplicative pseudo-polynomial dependence on the largest input weight; this is known as the “(polynomial) embedding technique”. Notably, this situation also concerns TSP: Björklund proved that Hamiltonicity can be leveraged to solve TSP by paying the price of a near-linear time-factor depending on the largest edge weight $W$ [Bjö10, Thm. 3]. This becomes impractical for arbitrarily large $W$. The same holds true for the weighted Max-Cut problem: its unweighted counterpart can be solved in $O^{*}(2^{\omega n/3})$-time [Wil05, Koi06], where $\omega<2.371339$ is the matrix multiplication exponent [ADV+25], while its weighted version has to “accept” the additional pseudo-polynomial factor. On the other hand, if the weights are polynomially bounded, then the problematic factor becomes polynomial, and the running time of the weighted version matches, in the $O^{*}$-sense, that of the unweighted version.
@@
Given the ubiquity of the embedding technique, it is natural to ask whether the gap between the running times of the weighted and unweighted versions of a problem is overcome only in the regime of polynomially bounded weights. A negative answer would be consistent with the widespread adoption of the embedding approach, with its inherent pseudo-polynomial dependence [Wil05, Koi06, BHKK07, BHK09, Bjö10, LN10, CFK+15]. In contrast, a positive answer would re-open the door to a more systematic re-examination of the weighted vs. unweighted gap, in the spirit of the “Losing Weights by Gaining Edges” technique for node-weighted problems [ALW14] or in terms of approximation guarantees [CST01].
@@ 1.1 Our Results
To maintain notation consistency with prior work on unweighted problems [RW24], we prefix problem names with $\mathcal{C}$-” if the set of input weights has doubling constant $\mathcal{C}$. We also write $O_{\mathcal{C}}$ to note that we suppressed factors that depend only on $\mathcal{C}$. Under this notation, all of our results conform to the following template: {restatable}[]corollaryDCTSP If Hamiltonicity can be solved in time $T(n)$, then $\mathcal{C}$-TSP can be solved in time $O_{\mathcal{C}}^{*}(T(n))$. The applicability of our work goes beyond TSP. Indeed, we can consider any weighted NP-hard problem that operates over the $(\min,+)$ and $(\max,+)$ semirings, such as the Weighted Max-Cut, Edge-Weighted $k$-Clique, or Minimum Steiner Tree problems. [^3] We kindly refer the reader to the relevant section for the problem definitions. The only requirement is that the problem at hand satisfies a property $\phi$: its objective value is an additive combination of input weights, and it admits an (algebraic) algorithm as seen earlier; see Prop. 1 for a formal treatment.
@@
[Meta-algorithm]theoremMETATHM If problem $\mathbf{P}_{w}$ satisfies property $\phi$ and the unweighted version can be solved in time $O(T(n))$ by an algebraic algorithm $\mathcal{A}$, then $\mathcal{C}$-$\mathbf{P}_{w}$ can be solved in time $O_{\mathcal{C}}^{*}(T(n))$.
@@ 1.2 Overview of Approach
Let $A$ denote the set of input weights of the problem at hand, such as the edge weights in TSP. The first observation is that the objective value of the problem is contained in a $q$-fold sumset of $A$, i.e., $A+\ldots+A$ with $A$ occurring $q$ times, where $q$ is problem-specific; for TSP we have $q=n$. The second observation is that if $A$ has small doubling, then the cardinality of this very $q$-fold sumset is polynomial w.r.t. $O_{\mathcal{C}}$. [^4] Noteworthy, Plünnecke’s inequality is too loose in this setting, leading to an exponential bound. This suggests that, in principle, we could run the bounded-input algorithm. However, the values may still be numerically as large as $\Theta(q\cdot\max A)$. The missing piece is a decomposition of the weights into constant-size coefficient-tuples coming from the generalized arithmetic progression $A$ is a subset of (cf. Freiman’s theorem [Fre73]).
@@ 1.3 Related Work
Our work continues the effort to consider important results from additive combinatorics [TV06] as a new toolbox for algorithm design, a connection highlighted in works such as Ref. [Tre09, Vio11, Bib13, Lov17]. One of the first seminal results is that of Chan and Lewenstein [CL15], who used the (constructive) Balog-Szemerédi-Gowers (BSG) theorem to speed up the monotone variant of the min-plus convolution on small integers. Recently, this has been further used and extended in #APSP [CVX23], 3SUM [JX23], listing 4-cycles [ABF23], and approximate distance oracles [ABF23]. We observe that the constructive version of a fundamental theorem in additive combinatorics bears fruit in algorithm design. To this end, Randolph and Węgrzycki [RW24] made Freiman’s theorem [Fre73], another fundamental result in additive combinatorics, constructive. Their main applications were to SubsetSum and ILP-Feasibility, which have an inherent additive structure. Indeed, we extend their results to the area of weighted problems.
@@
Repurposing the unweighted algorithm for the weighted version of a problem is a common theme in work that improves over the textbook solution, as in seminal results such as Ref. [Wil05, BHKK07, BHK09, Bjö10]; notably, in all these examples, the weighted problem operates over the $(\min,+)$ and $(\max,+)$ semirings. The key idea is to represent the input weights of the problem instance as monomials so that the bounded-input algorithm can work in the $(+,\times)$ ring. The drawback is that this introduces a pseudo-polynomial dependence in the running time, since operations now incur an additional $\widetilde{O}(W)$ factor, [^5] The $\widetilde{O}-$notation hides polylogarithmic factors in the input size. where $W$ is the largest input weight. For instance, undirected Hamiltonicity runs in $O^{*}(1.66^{n})$-time, hence TSP can be solved in $O^{*}(1.66^{n}W)$-time. This, however, becomes impractical for arbitrarily large $W$. Still, if $W$ is polynomially bounded, then up to polynomial factors, the weighted problem becomes equivalent to the unweighted one. For example, if $W=\text{poly}(n)$, then TSP can be solved in time $O^{*}(1.66^{n})$, matching its unweighted counterpart.
@@
The pseudo-polynomial factor is often unwanted. One way to shave it is to enter the realm of approximation. In fact, there is a simple exercise to turn any bounded-input algorithm for a problem into an exponential-time $(1+\varepsilon)$-approximation algorithm. For the Max-Cut problem, Williams mentions this idea in passing [Wil05], inspired by the literature on the all-pairs shortest paths problem [Zwi02]; see, e.g., Ref. [Sto24] for more such examples. It is interesting to ask whether the doubling constant could play a role in approximation algorithms.
@@ 2 Preliminaries
As we will only deal with exponential-time algorithms, we use the $O^{*}$-notation to hide polynomial factors in the input size. For example, $O^{*}(2^{n})=O(2^{n}\text{poly}(n))$. In particular, this notation will never hide pseudo-polynomial factors depending on the input weights; specifically, we mean the largest input weight, which will be denoted by $W$. Furthermore, we use the $O_{\mathcal{C}}$-notation, already established in Ref. [RW24], to denote that we suppressed factors depending on $\mathcal{C}$, the doubling constant on which we parameterize the time-complexities. For instance, $O_{\mathcal{C}}(2^{n})=f(\mathcal{C})\cdot O(2^{n})$ for a computable function $f$. To put into context, $O_{\mathcal{C}}^{*}(2^{n})$ will thus mean $f(\mathcal{C})\cdot O^{*}(2^{n})$.
@@ 2.1 Embedding Technique
The solution polynomial abstracts away whether the problem is a minimization or maximization problem: for minimization we select the monomial with the smallest exponent, whereas for maximization we select the monomial with the largest exponent. To this end, we refer to an algorithm as algebraic, if it returns a solution polynomial for an instance of a problem. [^7] An alternative attribute could be non-combinatorial. For a treatment on the definition of a combinatorial algorithm, see Ref. [AFK+24].
@@ 2.2 Additive Combinatorics
Given a set $A$ of small doubling, Freiman’s theorem [Fre73] tells that, indeed, $A$ is included in a GAP $G$, whose dimension is independent on $|A|$. In this work, we will need the constructive Freiman’s theorem, recently designed by Randolph and Węgrzycki [RW24]:
Theorem 2.2 (Constructive Freiman’s Theorem [RW24]).
@@
The above theorem gives us a GAP of dimension $d(\mathcal{C})=2^{\mathcal{C}^{O(1)}}$ and volume $v(\mathcal{C})=2^{2^{\mathcal{C}^{O(1)}}}n$, as in the original Freiman’s theorem [Fre73]. Noteworthy, later works do optimize $d$ and $v$ [Cha02, Sch11, San12a, San12b], yet these improvements are not captured by the constructive version.
@@ 4 Applications
In the following, we review several weighted NP-hard problems and show how to solve their small-doubling instances faster than the traditional algorithm by leveraging the previous meta-algorithm. This enables running times in the regime of the unweighted version of the problem for a large suite of problems, such as TSP, Weighted Max-Cut, Edge-Weighted $k$-Clique, and Minimum Steiner Tree. We stress that, compared to prior work [RW24], we consider weighted problems that operate over tropical semirings.
@@ 4.1 TSP Meets Small Doubling
Condition 1 holds since a tour’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as Björklund’s algorithm [Bjö10] for undirected Hamiltonicity, which runs in time $O^{*}(1.66^{n})$, and solves the weighted case in time $O^{*}(1.66^{n}\,W)$ [Bjö10, Thm. 3]. Due to the algorithm’s design, namely that it counts cycle covers of $n$ arcs, we have $\lambda\leq n$. ∎
@@ 4.2 Weighted Max-Cut Under Small Doubling
Another well-studied problem whose unweighted version has been sped up and fits the small doubling setting well is the Weighted Max-Cut problem. In a breakthrough result, Williams [Wil05] improved the time complexity of the unweighted Max-Cut to $O^{*}(2^{\omega n/3})$. We show that we can leverage this result to the weighted setting when the edge weights form an integer set that has small doubling. We formally define the doubling constant parametrization of the problem in the following:
@@
Condition 1 holds since a cut’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as Williams’s algebraic algorithm [Wil05] (and later improved by Koivisto [Koi06]) for the unweighted Max-Cut problem, which runs in time $O^{*}(2^{\omega n/3})$, and solves the weighted case in time $O^{*}(2^{\omega n/3}\,W)$ [Koi06, Thm. 3]. ∎
@@ 4.3 Small Doubling and Edge-Weighted $k$-Cliques
We now turn to a problem whose complexity is a rather popular hypothesis in hardness proofs [BT17, AVW14, LVW18]: the edge-weighted $k$-clique problem. Compared to the problems considered so far, the algebraic algorithm for the unweighted $k$-Clique has already been known from the 80’s [NP85]. We show that it can be leveraged in the doubling constant parametrization of the edge-weighted version. Note that, apart from a mention in passing in Lincoln et al. [LVW18] that Nešetřil-Poljak’s algorithm can also be used for polynomially bounded weights, we could not find any reference to a proof. For completeness, we show this in Lemma B (Appendix B).
@@
Condition 1 holds since a clique’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as Nešetřil-Poljak’s algebraic algorithm [NP85] for the unweighted version of the problem (or the improvement by Eisenbrand and Grandoni [EG04] for some values of $k$ using rectangular matrix multiplication), which runs in time $O^{*}(n^{k\omega/3})$, and solves the weighted case in time $O^{*}(n^{k\omega/3}\,W)$ (Lemma B). Notably, due to its design, namely that each original input weight is doubled in the auxiliary graph, the bounded-input algorithm (Lemma B) has a $\lambda\leq 2|w(I)|$. ∎